The least perfect square, which is divisible by each of 2136 & 66 is:

Grade:10

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

if we look that 2136 * 66 = 89 *22*62*11
So, the least perfect square that would be divisible each by 2136 and 66 would be equal to 892*22* 62*112= 138015504

thanks